OpenStudy (anonymous):

given a function 'g' which has a derivative g'(x) for every real 'x' and which satisfy g'(0)=2 and g(x+y)=e^y*g(x) + e^x*g(y) foe all x & y. Find g(x) pl. help immediate medals awarded

OpenStudy (anonymous):

maybe this would help f'(x)=lim h->0 [f(x+h)-f(x)]/h

OpenStudy (anonymous):

\[g(x+y)=e^yg(x) + e^xg(y)\] Differentiate \[g'(x+y) \times (1+ \frac{dy}{dx})=e^y g'(x)+e^yg(x) \frac{dy}{dx}+e^x g'(y) \frac{dy}{dx}+e^x g(y) \}\] \[\frac{dy}{dx}=0\] \[g'(x+y)=e^y g'(x)+e^x g(y)\] Replace y with x and x with 0 \[g'(x)=e^x g'(0)+e^0 g(x)\] \\[\LARGE g(x)=g'(x)-2\] Replace x with 0 \[g(0)=0\] hmm..gotta think about rest,just tried :|

OpenStudy (anonymous):

can there be a more simplified answer?curious :o

OpenStudy (anonymous):

@rishabhjaiswal225499 what is the answer?

OpenStudy (anonymous):

the answer 2x(e^(2x))

OpenStudy (anonymous):

did u get what i did?

OpenStudy (anonymous):

yup i got but couldt not do further

OpenStudy (anonymous):

Differentiate again \[g''(x)=g'(x)\] Do you know any function whose differential is it itself?

OpenStudy (anonymous):

It is \(ce^x\)

OpenStudy (anonymous):

ok got it now

OpenStudy (anonymous):

thanks